# Chromatic numbers, Buchstaber numbers and chordality of Bier spheres

### v1

@@ 3. Bier spheres and their chromatic numbers
Since $\chi(\mathrm{Bier}(K))=m$ and $\mathrm{Bier}(K)$ is a suspension over $\mathrm{Bier}(R)$, by Lemma 3.7 we get:
@@
    $\chi(\mathrm{Bier}(K))=m-1$ if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next two cases holds:
    • $P_{K}=I^{m-1}$;
    • $P_{K}=I^{m-3}\times P_{6}$.
@@
The case $m=2$ is clear. In case $m\geq 3$, statement (a) follows directly from Proposition 3.1, while statement (b) is a direct consequence of Lemma 3.15. ∎
@@ 4. Buchstaber numbers and chordality of Bier spheres
Note that this definition is equivalent to the one given in [6, Definition 5.3]; moreover, in the case $p=2$, it is equivalent to the one of $s_{\mathbb{R}}(K)$ as the maximal rank of a real toric subgroup in $\mathbb{Z}_{2}^{m}$ acting freely on the real moment-angle-complex $\mathcal{R}_{K}$ and similarly in the complex case, see [6, Definition 5.1, Remark 5.5].
It follows from the above definition that both $s(K)$ and $s_{p}(K)$, for any prime $p\in\mathbb{N}$, are combinatorial invariants of $K$. The real Buchstaber number was introduced and studied in [14], where it was proven that
$1\leq s(K)\leq s_{\mathbb{R}}(K)\leq m-n,$
for any simplicial complex $K$ with $f_{0}(K)=m$ and $\dim(K)=n-1$, see also [1, Proposition 3] for an alternative proof. The mod $p$ Buchstaber invariant was introduced and studied in [6], where it was shown, see [6, Corollary 5.8], that
$m-\chi(K)\leq s(K)\leq s_{p}(K)\leq m-n,\text{ for any prime $p\in\mathbb{N}$}.$
Since $\dim(\mathrm{Bier}(K))=m-2$, applying these inequalities and Theorem 3.16 to a Bier sphere, one has:
$f_{0}(\mathrm{Bier}(K))-m\leq s(\mathrm{Bier}(K))\leq s_{p}(\mathrm{Bier}(K))\leq f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
@@
In [21, Theorem 3.4], the real and complex Buchstaber invariants of $\mathrm{Bier}(K)$ were computed assuming that $K$ is considered alongside with its ghost vertices. Now, we are going to state and prove the general formulae for all Buchstaber numbers of $\mathrm{Bier}(K)$, when only geometrical vertices of $K$ are taken into account.
@@
$s(\mathrm{Bier}(K))=s_{p}(\mathrm{Bier}(K))=f_{0}(K)-f_{m-2}(K)+1,\text{ for any prime }p\in\mathbb{N}.$
@@
First, let us compute the number of geometrical vertices of $\mathrm{Bier}(K)$ in terms of the $f$-vector of $K$. By [21, Corollary 2.7], the Bier sphere $\mathrm{Bier}(K)$ has $(m-f_{0}(K))+f_{m-2}(K)$ ghost vertices, which is just the sum of the numbers of ghost vertices in $K$ and in its Alexander dual complex $K^{\vee}$. Hence the number of geometrical vertices of the Bier sphere equals
@@
$\phi\colon[m]\sqcup[m^{\prime}]\to\mathbb{Z}^{m-1}\text{ such that }\phi\colon i,i^{\prime}\mapsto e_{i},\text{ for all }1\leq i\leq m-1,\text{ and }\phi\colon m,m^{\prime}\mapsto e_{1}+e_{2}+\ldots+e_{m-1}.$
Let us define $\Lambda_{(p)}$ to be the restriction of $\phi$ to $V(\mathrm{Bier}(K))$. Then $\Lambda_{(p)}$ is a characteristic map for $\mathrm{Bier}(K)$ both in the complex and mod $p$ case, for each prime $p\in\mathbb{N}$. This follows from the fact that maximal simplices of $\mathrm{Bier}(K)$ have cardinality $(m-1)$ and do not contain a pair $\{i,i^{\prime}\}$, for any $1\leq i\leq m$.
@@
$s_{p}(K)\geq s(K)\geq r=f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
@@
$s(K)\leq s_{p}(K)\leq f_{0}(\mathrm{Bier}(K))-(\dim(\mathrm{Bier}(K))+1)=f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
Thus one gets the following equality
$s(K)=s_{p}(K)=f_{0}(\mathrm{Bier}(K))-(m-1),\text{ for any prime }p\in\mathbb{N}.$
Finally, the right hand side equals $(m+f_{0}(K)-f_{m-2}(K))-(m-1)=f_{0}(K)-f_{m-2}(K)+1$. Q.E.D. ∎
@@
We write $P=\mathrm{vc}^{k}(\Delta^{n})$ with $n\geq 2,k\geq 0$ and say that $P$ is a truncation polytope if $P$ is combinatorially equivalent to a convex simple $n$-dimensional polytope obtained from $\Delta^{n}$ by a sequence of $k$ vertex truncations by means of hyperplanes in general position.
@@
Note that the polytope $Q$ combinatorially dual to a truncation polytope $P$ is combinatorially equivalent to a convex simplicial polytope obtained from $\Delta^{n}$ by a sequence of $k$ stellar subdivisions in facets. Such a polytope is called a stacked polytope. Observe that the combinatorial types of $P$ and $Q$ depend on the vertices truncated and facets subdivided, respectively, when $n\geq 3$ and $k\geq 3$. Namely, there is only one combinatorial type of $\mathrm{vc}^{0}(\Delta^{3})$, $\mathrm{vc}^{1}(\Delta^{3})$, and $\mathrm{vc}^{2}(\Delta^{3})$, whereas there are 3 combinatorially different realizations of $\mathrm{vc}^{3}(\Delta^{3})$ (see Figure 7) and 7 combinatorially different realizations of $\mathrm{vc}^{4}(\Delta^{3})$ [16, Table 8].
@@
Let $P=\mathrm{vc}^{k}(\Delta^{n})$ with $n\geq 3,k\geq 0$ be a truncation polytope. Then $K_{P}$ is chordal.
@@
To make the induction step, suppose $P=\mathrm{vc}^{k+1}(\Delta^{n})$ and $n\geq 3,k\geq 0$. Then the dual stacked polytope $Q$ was obtained by performing a stellar subdivision with apex $v$ in a facet $\sigma:=\{i_{1},\ldots,i_{n}\}$ of a stacked polytope $Q^{\prime}$ which is dual to a certain simple polytope of the type $\mathrm{vc}^{k}(\Delta^{n})$.
@@
If $v\notin S$, then $S$ was an induced cycle in $\partial Q^{\prime}$, which is chordal by inductive assumption, and hence $p=3$, a contradiction. Indeed, since $\sigma$ was a simplex in $\partial Q^{\prime}$, the induced chordless cycle $S$ of length $p\geq 4$ could intersect it in no more than two vertices. Then $S$ remains a full subcomplex in $\partial Q$ after the stellar subdivision in $\sigma$ was performed.
@@
We are ready to state the result, which provides the classification of all chordal Bier spheres in dimensions greater than one.
@@
The implication $\Leftarrow$ follows immediately from Proposition 4.5. Indeed, the case of a void complex is obvious. Suppose $k>0$ and $K^{\vee}$ is generated by its vertex set $\{1^{\prime},2^{\prime},\ldots,k^{\prime}\}$. By definition of a Bier sphere, it implies that $\mathrm{Bier}(K)$ is obtained from $\partial\Delta_{[m]}$ by stellar subdivisions of the facets $[m]\setminus\{i\}$ by vertices $i^{\prime}$ for $1\leq i\leq k$, see Figure 6. We obtain a polytopal Bier sphere, whose Bier polytope $P_{K}$ is of the type $\mathrm{vc}^{k}(\Delta^{m-1})$.
@@
To prove the implication $\Rightarrow$, suppose that $\mathrm{Bier}(K)$ is chordal and $\mathrm{sk}^{1}(K)$ contains an edge. Without loss of generality, we can assume that $\{1,2\}\in K$.
@@
Then $\{1^{\prime},2^{\prime}\}\notin K^{\vee}$, otherwise $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,1^{\prime},2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. It follows that if $\{i\}\notin K$ with $i\geq 3$, then Alexander duality would imply that $\{1^{\prime},2^{\prime}\}\subset[m^{\prime}]\setminus\{i^{\prime}\}\in K^{\vee}$, which contradicts $\{1^{\prime},2^{\prime}\}\notin K^{\vee}$. This yields that $f_{0}(K)=m$ and $K$ has no ghost vertices.
@@
Now, we are going to show that both $\{1,i\}$ and $\{2,i\}$ are in $K$, for each $i$ with $3\leq i\leq m$. Indeed, otherwise one of the following three cases occurs. If there exists an $i,3\leq i\leq m$ such that $\{1,i\},\{2,i\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 5-cycle on the vertices $\{1,2,1^{\prime},i,2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. If there exists an $i,3\leq i\leq m$ such that $\{1,i\}\notin K$ and $\{2,i\}\in K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,i,2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. Finally, if there exists an $i,3\leq i\leq m$ such that $\{1,i\}\in K$ and $\{2,i\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,1^{\prime},i\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction.
@@
Next, observe that $\{i,j\}\in K$, for every pair of vertices $i,j$ such that $3\leq i\neq j\leq m$. Indeed, if $\{i,j\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,i,1^{\prime},j\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction.
@@
Finally, it follows that $K^{\vee}$ contains no edges. Indeed, as was already shown above, $\mathrm{sk}^{1}(K)$ is a complete graph with $m$ vertices. Suppose there was an edge $\{i^{\prime},j^{\prime}\}\in K^{\vee}$ with $1\leq i\neq j\leq m$. Then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{i,j,i^{\prime},j^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. This finishes the proof. ∎
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then $\mathrm{Bier}(K)$ is chordal if and only if $\mathrm{Bier}(K)$ is polytopal and one of the next three cases holds:
@@
  • $P_{K}=\mathrm{vc}^{k}(\Delta^{m-1})$, where $0\leq k\leq m$ and $m\geq 4$.
@@
In case $m=2$ the unique 00-dimensional Bier sphere is a chordal complex. In case $m=3$ the only chordal $1$-dimensional Bier sphere is the boundary of a triangle, see Figure 1. If $m\geq 4$, then the application of Lemma 4.6 finishes the proof. ∎

### v2

@@ 3. Bier spheres and their chromatic numbers
Since $\chi(\mathrm{Bier}(K))=m$ and $\mathrm{Bier}(K)$ is a suspension over $\mathrm{Bier}(R)$, by Lemma 3.7, we get:
@@
    $\chi(\mathrm{Bier}(K))=m-1$ if and only if $\mathrm{Bier}(K)$ is polytopal and the next equivalent conditions hold:
    • $K\in\{C^{m-3}(\Gamma_{4})$, $C^{m-3}(\Gamma_{4}^{\vee})$, $C^{m-3}(G_{4})$, $C^{m-3}(\Gamma_{6})\}$;
    • $P_{K}\in\{I^{m-1},I^{m-3}\times P_{6}\}$.
@@
The case $m=2$ is clear. In case $m\geq 3$, statement (a) follows directly from Proposition 3.1, while statement (b) is a direct consequence of Lemma 3.15 and the fact that the operations of taking a cone of a simplicial complex and taking its Alexander dual commute. ∎
@@ 4. Buchstaber numbers and chordality of Bier spheres
Note that this definition is equivalent to the one given in [6, Definition 5.3]; moreover, in the case $p=2$, it is equivalent to the one of $s_{\mathbb{R}}(K)$ as the maximal rank of a real toric subgroup in $\mathbb{Z}_{2}^{m}$ acting freely on the real moment-angle-complex $\mathcal{R}_{K}$ and similarly in the complex case, see [1, 16, 17, 18].
It follows from the above definition that both $s(K)$ and $s_{p}(K)$, for any prime $p\in\mathbb{N}$, are combinatorial invariants of $K$. Buchstaber invariant of a simplicial complex is known to be closely related to its chromatic number. First, given a simplicial complex $K$ with $f_{0}(K)=m$ and $\dim(K)=n-1$, the fundamental inequality $s(K)\leq m-n$ is well-known and can be found in [22]. In the same paper, the Buchstaber and chromatic numbers were linked by the inequality $m-\chi(K)\leq s(K)$. Furthermore, for any simplicial complex, the real Buchstaber number bounds the complex Buchstaber number from above: $s(K)\leq s_{2}(K)$, see [16, Fact 15, p.5] as well as [1, 18, 17]. Finally, for every prime $p\neq 2$, this is proved by literally the same argument, see [6], where the mod $p$ Buchstaber invariant was introduced and studied, and therefore the chain of inequalities takes place:
$m-\chi(K)\leq s(K)\leq s_{p}(K)\leq m-n\text{ for any prime $p\in\mathbb{N}$}.$
Since $\dim(\mathrm{Bier}(K))=m-2$, applying these inequalities and Theorem 3.16 to a Bier sphere, we get:
$f_{0}(\mathrm{Bier}(K))-m\leq s(\mathrm{Bier}(K))\leq s_{p}(\mathrm{Bier}(K))\leq f_{0}(\mathrm{Bier}(K))-(m-1)\text{ for any prime }p\in\mathbb{N}.$
@@
In [27, Theorem 3.4], the real and complex Buchstaber invariants of $\mathrm{Bier}(K)$ were computed assuming that $K$ is considered alongside with its ghost vertices. Now, we are going to state and prove the general formulae for all Buchstaber numbers of $\mathrm{Bier}(K)$, when only geometrical vertices of $K$ are taken into account.
@@
[…]
@@
Suppose $\mathrm{Bier}(K)$ is a polytopal sphere. Recall that by Theorem 3.16 (a), for any $\mathrm{Bier}(K)$ with $K$ on $m],m\geq 2$ one has either $\chi(\mathrm{Bier}(K))=m-1$, or $\chi(\mathrm{Bier}(K))=m$. Note that when $\chi(\mathrm{Bier}(K))=m-1=\dim(\mathrm{Bier}(K))+1$, the above inequalities immediately imply that the complex and all mod $p$ Buchstaber numbers of a Bier sphere are maximal possible in the first case. When $\chi(\mathrm{Bier}(K))=m=\dim(\mathrm{Bier}(K))+2$, it follows from [15, Theorem (7)] that the chain of equalities and inequalities
$s(\mathrm{Bier}(K))\geq f_{0}(\mathrm{Bier}(K))-\chi(\mathrm{Bier}(K))+s(\Delta^{\chi(\mathrm{Bier}(K))-1}_{\dim(\mathrm{Bier}(K))})=\\ =f_{0}(\mathrm{Bier}(K))-m+s(\Delta^{m-1}_{m-2})\geq f_{0}(\mathrm{Bier}(K))-m+1$
holds, where $\Delta_{p}^{q}$ denotes the $p$-skeleton of the $q$-dimensional simplex. This shows that the complex and all mod $p$ Buchstaber numbers of a Bier sphere are also maximal possible in the second case. Now we consider the case of a general $\mathrm{Bier}(K)$ and we are going to introduce a canonical characteristic map for every Bier sphere. First, let us compute the number of geometrical vertices of $\mathrm{Bier}(K)$ in terms of the $f$-vector of $K$. By [27, Corollary 2.7], the Bier sphere $\mathrm{Bier}(K)$ has $(m-f_{0}(K))+f_{m-2}(K)$ ghost vertices, which is just the sum of the numbers of ghost vertices in $K$ and in its Alexander dual complex $K^{\vee}$. Hence, the number of geometrical vertices of the Bier sphere equals
@@
[…]
Let us define $\Lambda_{(p)}$ to be the restriction of $\phi$ to $V(\mathrm{Bier}(K))$. Then $\Lambda_{(p)}$ is a characteristic map for $\mathrm{Bier}(K)$, both in the complex and mod $p$ case for each prime $p\in\mathbb{N}$. This follows from the fact that maximal simplices of $\mathrm{Bier}(K)$ have cardinality (m1)(m-1) and do not contain a pair $\{i,i^{\prime}\}$ for any $1\leq i\leq m$.
@@
$s_{p}(K)\geq s(K)\geq r=f_{0}(\mathrm{Bier}(K))-(m-1)\text{ for any prime }p\in\mathbb{N}.$
@@
[…]
Thus, one gets the following equality:
[…]
Finally, the right hand side equals $(m+f_{0}(K)-f_{m-2}(K))-(m-1)=f_{0}(K)-f_{m-2}(K)+1$. ∎
@@
We write $P\in\mathrm{vc}^{k}(\Delta^{n})$ with $n\geq 2,k\geq 0$ and say that $P$ is a truncation polytope if $P$ is combinatorially equivalent to a convex simple $n$-dimensional polytope obtained from $\Delta^{n}$ by a sequence of $k$ vertex truncations by means of hyperplanes in general position.
@@
Note that the polytope $Q$ combinatorially dual to a truncation polytope $P$ is combinatorially equivalent to a convex simplicial $n$-dimensional polytope obtained from $\Delta^{n}$ by a sequence of $k$ stellar subdivisions in facets. Such a polytope is called a stacked polytope. Observe that the combinatorial types of $P$ and $Q$ depend on the vertices truncated and facets subdivided, respectively, when $n\geq 3$ and $k\geq 3$. Namely, there is only one combinatorial type of $\mathrm{vc}^{0}(\Delta^{3})$, $\mathrm{vc}^{1}(\Delta^{3})$, and $\mathrm{vc}^{2}(\Delta^{3})$, whereas there are 3 combinatorially different realizations of $\mathrm{vc}^{3}(\Delta^{3})$ (see Figure 7) and 7 combinatorially different realizations of $\mathrm{vc}^{4}(\Delta^{3})$ [20, Table 8].
@@
The next result follows from [26, Theorem 8.5]. We give it with proof here for the sake of completeness.
@@
Let $P\in\mathrm{vc}^{k}(\Delta^{n})$ with $n\geq 3,k\geq 0$ be a truncation polytope. Then $K_{P}$ is chordal.
@@
To make the induction step, suppose $P\in\mathrm{vc}^{k+1}(\Delta^{n})$ and $n\geq 3,k\geq 0$. Then the dual stacked polytope $Q$ was obtained by performing a stellar subdivision with apex $v$ in a facet $\sigma=\{i_{1},\ldots,i_{n}\}$ of a stacked polytope $Q^{\prime}$ which is dual to a certain simple polytope in $\mathrm{vc}^{k}(\Delta^{n})$.
@@
If $v\notin S$, then $S$ was an induced cycle in $\partial Q^{\prime}$, which is chordal by inductive assumption, hence $p=3$, a contradiction. Indeed, since $\sigma$ was a simplex in $\partial Q^{\prime}$, the induced chordless cycle $S$ of length $p\geq 4$ could intersect it in no more than two vertices. Then $S$ remains a full subcomplex in $\partial Q$ after the stellar subdivision in $\sigma$ was performed.
@@
The converse of the previous statement is not true; e.g. if $P$ was obtained from $\Delta^{2}\times\Delta^{2}$ by a vertex cut, then $P$ is not a truncation polytope, but $K_{P}$ is a chordal complex. We are ready to state the result, which provides the classification of all chordal Bier spheres in dimensions greater than one. It implies that for Bier spheres of dimension greater than one, being a stacked sphere and being a chordal complex are equivalent.
@@
The implication $\Leftarrow$ follows immediately from Proposition 4.5. Indeed, the case of a void complex is obvious. Suppose $k>0$ and $K^{\vee}$ is generated by its vertex set $\{1^{\prime},2^{\prime},\ldots,k^{\prime}\}$. It follows from the definition of a Bier sphere that $\mathrm{Bier}(K)$ is obtained from $\partial\Delta_{[m]}$ by stellar subdivisions of the facets $m]\setminus\{i\}$ by vertices $i^{\prime}$ for $1\leq i\leq k$. The corresponding simple polytope $P_{K}$ is then a truncation polytope obtained from $\Delta^{m-1}$ by cutting off $k$ of its $m$ vertices, see Figure 6 for the case $m=4$ and $0\leq k\leq 4$. Therefore, we obtain a polytopal Bier sphere, whose Bier polytope $P_{K}$ is of the type in $\mathrm{vc}^{k}(\Delta^{m-1})$.
@@
To prove the implication $\Rightarrow$, suppose that $\mathrm{Bier}(K)$ is chordal and each of the graphs, $\mathrm{sk}^{1}(K)$ and $\mathrm{sk}^{1}(K^{\vee})$, contains an edge. Without loss of generality, we can assume that $\{1,2\}\in K$.
@@
Then $\{1^{\prime},2^{\prime}\}\notin K^{\vee}$, otherwise $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,1^{\prime},2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. Note that if $\{i\}\notin K$ for some $i\geq 3$, then Alexander duality would imply that $\{1^{\prime},2^{\prime}\}\subset[m^{\prime}]\setminus\{i^{\prime}\}\in K^{\vee}$, which contradicts $\{1^{\prime},2^{\prime}\}\notin K^{\vee}$. It follows that $f_{0}(K)=m$ and $K$ has no ghost vertices. The same argument shows that $K^{\vee}$ also has no ghost vertices.
@@
Now, we are going to show that both $\{1,i\}$ and $\{2,i\}$ are in $K$ for each $i$ with $3\leq i\leq m$. Indeed, otherwise one of the following three cases occurs. If there exists an $i,3\leq i\leq m$ such that $\{1,i\},\{2,i\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 5-cycle on the vertices $\{1,2,1^{\prime},i,2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. If there exists an $i,3\leq i\leq m$ such that $\{1,i\}\notin K$ and $\{2,i\}\in K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,i,2^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. Finally, if there exists an $i,3\leq i\leq m$ such that $\{1,i\}\in K$ and $\{2,i\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,2,1^{\prime},i\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction.
@@
Next, observe that $\{i,j\}\in K$ for every pair of vertices $i,j$ such that $3\leq i\neq j\leq m$. Indeed, if $\{i,j\}\notin K$, then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{1,i,1^{\prime},j\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction.
@@
Thus, $\mathrm{sk}^{1}(K)$ is a complete graph with $m$ vertices. Suppose there was an edge $\{i^{\prime},j^{\prime}\}\in K^{\vee}$ with $1\leq i\neq j\leq m$. Then $\mathrm{Bier}(K)$ contains a chordless 4-cycle on the vertices $\{i,j,i^{\prime},j^{\prime}\}$ as a full subcomplex and therefore $\mathrm{Bier}(K)$ is non-chordal, a contradiction. It implies that $\mathrm{sk}^{1}(K^{\vee})$ has no edges, which contradicts our initial assumption. This finishes the proof. ∎
@@
Let $K\neq\Delta_{[m]}$ be a simplicial complex on $[m]$ with $m\geq 2$. Then the following three statements are equivalent:
  • (a)
    $\mathrm{Bier}(K)$ is a chordal complex;
  • (b)
    $\mathrm{Bier}(K)$ is a polytopal sphere and one of the next three cases holds:
    1. (1)
@@
    3. (3)
      $P_{K}\in\mathrm{vc}^{k}(\Delta^{m-1})$, where $0\leq k\leq m$ and $m\geq 4$.
  • (c)
    One of the next three cases holds:
    1. (1)
      $m=2$;
    2. (2)
      $m=3$ and either $K$ or $K^{\vee}$ equals $\partial\Delta^{2}$;
    3. (3)
      $m\geq 4$ and either $K$ or $K^{\vee}$ has no edges.
If $m\leq 3$, then $\mathrm{Bier}(K)$ is a stacked sphere being either the boundary of a simplex, or the $k$-cycle with $k=4,5,$ or $6$. If $m\geq 4$, then $P_{K}$ is obtained by cutting off $k$ vertices of the simplex $\Delta^{m-1}$, where $k$ is the number of vertices of the corresponding complex without edges; in this case, (a), (b) and (c) are equivalent to:
  • (d)
    $\mathrm{Bier}(K)$ is a stacked sphere.
@@
Let us prove the equivalence of (a) and (c). In case $m=2$ the unique 0-dimensional Bier sphere is a chordal complex. In case $m=3$ the only chordal $1$-dimensional Bier sphere is the boundary of a triangle, see Figure 1. If $m\geq 4$, then the application of Lemma 4.6 finishes the proof of the equivalence (a) $\Leftrightarrow$ (c). The equivalence (b) $\Leftrightarrow$ (c) follows directly from the definition of Bier sphere. Finally, if $m\leq 3$, then $\mathrm{Bier}(K)$ is a stacked sphere due to Example 2.3; for $m\geq 4$, the implication (b) $\Rightarrow$ (d) is obvious, while the implication (d) $\Rightarrow$ (a) follows from Proposition 4.5. ∎